ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 232, 2019

Aramayona, J. and C. Martínez-Pérez Representations of pure symmetric automorphism groups of RAAGs

351

Bachoc, C., A. Gundert and A. Passuello The theta number of simplicial complexes

443

Baruch, E. M. and S. Purkait Newforms of half-integral weight: the minus space of S_{k+1/2}(Γ_{0}(8M))

41

Barvinok, A. Computing permanents of complex diagonally dominant matrices and tensors

931

Belolipetsky, M. and A. Lubotzky Counting non-uniform lattices  201

Braga, B. M. and A. Swift Coarse embeddings into superstable spaces

1

Bruin, H., D. Terhesiu and M. Todd The pressure function for infinite equilibrium measures

775

Bu, S. and G. Cai Periodic solutions of fractional degenerate differential equations with delay in Banach spaces

695

Burago, D., S. Ivanov and Y. Kurylev Spectral stability of metric-measure Laplacians

125

Cai, G. See S. Bu

Chen, L. Surjective homomorphisms between surface braid groups  483

Chodounský, D. and O. Guzmán There are no P-points in Silver extensions

759

Climenhaga, V. and V. Cyr Positive entropy equilibrium states  899

Cyr, V. See V. Climenhaga Da Silva, A. and C. Kawan Lyapunov exponents and partial hyper-

bolicity of chain control sets on flag manifolds

947

Dokuchaev, M. and N. Sambonet Schur’s theory for partial projective representations

373

Elazar, B. Quasi-Nash varieties and Schwartz functions on them  589

Erchenko, A. and A. Katok Flexibility of entropies for surfaces of negative curvature

631

Fernández-Alcober, G. A. and I. de las Heras Commutators in finite p-groups with 2-generator derived subgroup

109

Fumagalli, F., F. Leinen and O. Puglisi A reduction theorem for nonsolvable finite groups

231

Gundert, A. See C. Bachoc Guzmán, O. See D. Chodounský Halevi, Y. and A. Hasson Strongly dependent ordered abelian

groups and Henselian fields

719

Harbater, D., J. Hartmann, D. Krashen, R. Parimala and V. Suresh Local-global Galois theory of arithmetic function fields

849

Hartmann, J. See D. Harbater Hasson, A. See Y. Halevi de las Heras, I. See G. A. Fernández-Alcober Ivanov, S. See D. Burago Kamae, T. See X. Wang Katok, A. See A. Erchenko Kawan, C. See A. Da Silva Kohlenbach, U., G. López-Acedo and A. Nicolae Moduli of

regularity and rates of convergence for Fejér monotone sequences

261

Krashen, D. See D. Harbater Kurylev, Y. See D. Burago Lechner, R. Dimension dependence of factorization problems:

Hardy spaces and SL_{n}^{∞}

677

Leinen, F. See F. Fumagalli López-Acedo, G. See U. Kohlenbach Lubotzky, A. See M. Belolipetsky Maleva, O. and D. Preiss Cone unrectifiable sets and

non-differentiability of Lipschitz functions

75

Martínez-Pérez, C. See J. Aramayona Nadler, D. and Z. Yun Spectral action in Betti Geometric Lang-

lands

299

Nicolae, A. See U. Kohlenbach Parimala, R. See D. Harbater Passuello, A. See C. Bachoc Preiss, D. See O. Maleva Puglisi, O. See F. Fumagalli Purkait, S. See E. M. Baruch

Rudnick, Z. and E. Waxman Angles of Gaussian primes  159

Russo, F. G. Embeddings of n-Engel marginal subgroups in compact groups

921

Sambonet, N. See M. Dokuchaev

Shaul, L. Completion and torsion over commutative DG rings  531
Sjöberg, H. and G. M. Ziegler Semi-algebraic sets of f-vectors  827

Sobota, D. and L. Zdomskyy Convergence of measures in forcing extensions

501

Soo, T. Finitary isomorphisms of some infinite entropy Bernoulli flows  883

Suresh, V. See D. Harbater Swift, A. See B. M. Braga Tacy, M. L^{p} estimates for joint quasimodes of semiclassical pseudo-

differential operators

401

Terhesiu, D. See H. Bruin Todd, M. See H. Bruin

Wang, X. and T. Kamae Selection rules preserving normality  427

Waxman, E. See Z. Rudnick Yun, Z. See D. Nadler Zdomskyy, L. See D. Sobota Ziegler, G. M. See H. Sjöberg